# Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) is a popular approach for enhancing Large Language Models (LLMs) by addressing their limitations in verifying facts and answering knowledge-intensive questions. As the research in LLM extends their capability to handle input modality other than text, e.g. image, several multimodal RAG benchmarks are proposed. Nonetheless, they mainly use textual knowledge bases as the primary source of evidences for augmentation. There still lack benchmarks designed to evaluate images as augmentation in RAG systems and how they leverage visual knowledge. We propose Visual-RAG, a novel Question Answering benchmark that emphasizes visual knowledge intensive questions. Unlike prior works relying on text-based evidence, Visual-RAG necessitates text-to-image retrieval and integration of relevant clue images to extract visual knowledge as evidence. With Visual-RAG, we evaluate 5 open-sourced and 3 proprietary Multimodal LLMs (MLLMs), revealing that images can serve as good evidence in RAG; however, even the SoTA models struggle with effectively extracting and utilizing visual knowledge [^1] Benchmark and evaluation codes will be available at https://github.com/LuciusLan/Visual-RAG..
Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries
Yin Wu1, Quanyu Long1, Jing Li2, Jianfei Yu3, Wenya Wang1 1 Nanyang Technological University, 2 Harbin Institute of Technology (Shenzhen), 3 Nanjing University of Science and Technology wuyi0023@e.ntu.edu.sg
@@ 1 Introduction
Figure 1: Comparison on query format and QA pipeline between InfoSeek-style and our Visual-RAG. InfoSeek-style: The entity in the query image must first be recognized to extract textual evidence, which is then used independently to answer the question. Visual-RAG: The query is text-only, but direct answers are rarely available in textual form. The model must retrieve relevant images and interpret them to generate an answer.
Benchmark Query Modality Knowledge Modality Retrieval Modality InfoSeek (Chen et al., 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] E-VQA (Mensink et al., 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] WebQA (Chang et al., 2022) [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] MRAG-Bench (Hu et al., 2024) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Visual-RAG (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Table 1: Comparing modalities of knowledge intensive VQA benchmarks. The link icon [Uncaptioned image] denotes that the two objects are paired, retrieving one will automatically link to the other. For InfoSeek and E-VQA, Yan and Xie (2024) demonstrated that by image-to-image retrieval on Wikipedia images, the Wikipedia articles containing those images are also retrieved at a significantly higher recall comparing to image-to-text retrieval. Similarly for WebQA, a question-to-caption (text-to-text) retrieval achieves better performance than text-to-image retrieval, as shown in Table 5 in Appendix. Our Visual-RAG enforces text-to-image retrieval for visual knowledge.
We have observed rapid advancement of Large Language Models (LLMs) in recent years. However, they are prone to generating hallucinated responses Ji et al. (2023). Retrieval-Augmented Generation (RAG) has emerged as a widely adopted approach to mitigate this challenge (Lewis et al., 2020; Karpukhin et al., 2020; Izacard and Grave, 2021). By retrieving and integrating relevant textual documents from external knowledge sources, it enables LLMs to answer knowledge-intensive questions more accurately. Yet, many real-world scenarios demand multimodal knowledge (e.g., visual cues or geographical maps), suggesting that text-only retrieval may be insufficient.
Building on the success of text-only RAG, various evaluation benchmarks have been proposed to assess the capability of Multimodal LLMs (MLLMs, a.k.a. Large Vision Language Models LVLM) in using multimodal RAG to address complex, multimodal knowledge intensive questions, including InfoSeek (Chen et al., 2023), Encyclopedic-VQA (Mensink et al., 2023), etc. They feature mixed-modal queries where the question is asked based on a paired image.
However, existing datasets reveal several limitations. Firstly, although images are included in the query, MLLMs do not utilize the visual knowledge from these images when generating answers. Secondly, questions in these datasets typically do not require reasoning over knowledge extracted from images. Figure 1 gives an example on InfoSeek-styled query. To answer the query, a model needs to first recognize the entity in query image, then locate the Wikipedia article and augment relevant passage to LLMs to generate an answer. Query image information is only required in the entity recognition step. These existing multimodal benchmarks are textual knowledge intensive, highlighting the need for a new benchmark for visual knowledge intensive QA.
To this end, We present Visual-RAG, a visual knowledge intensive QA benchmark designed to evaluate text-to-image retrieval and answer generation augmented by retrieved images. Key features of Visual-RAG are:
@@
    The entities in queries are obscure, with their descriptions being brief, rarely documented in standard references (e.g., Wikipedia)
@@
    The visual features queried are fine-grained, with only 5.35% of the images in the corpus containing the specific visual evidence needed to answer the questions, posing a challenging text-to-image retrieval task.
@@
    The majority of non-clue images are hard-negatives which depict the same entities and are visually similar. This subtle distinction further increases the difficulty of cross-modal retrieval.
To construct Visual-RAG, we begin with LLM-generated candidate queries, which are then refined through human filtering and rewriting. Next, we employ open-sourced MLLM and human annotators to perform another filtering step to ensure an appropriate level of difficulty for the queries. Detailed procedures for the entire benchmark construction can be found in Section 4.
With Visual-RAG, we comprehensively evaluate 8 mainstream MLLMs, including 5 open-sourced models and 3 proprietary ones. Firstly, we demonstrate that images can serve as powerful evidence for augmented generation through baseline experiments; secondly, we experiment with various RAG settings, testing the ability of MLLMs digesting images as augmentation, and analyse model behaviour on identifying clue from irrelevant images. Our benchmark sheds light on real-world scenarios for evaluating MLLMs, when users raise natural language queries for knowledge-intensive questions where textual evidence is scarce but visual evidence possibly exists, posing the challenging task of retrieving and utilizing relevant images in the wild.
The key findings from our experiments are summarized as follows:
@@
  • Cross-modal retrieval is challenging. The commonly used small cross-modal retriever (CLIP) faces significant challenges addressing difficult queries requiring the identification of fine-grained visual features.
@@
  • The amount of evidence affects retrieval-augmented generation. The evaluated MLLMs demonstrate the ability to extract visual knowledge to answer questions. However, a single ground-truth clue image is insufficient for the models to achieve optimal performance.
  • Negatives can be helpful. Providing multiple images as input, including both clue images and non-clue images, could enhance performance for proprietary models. This improvement likely stems from the models’ ability to contrast clue and non-clue images, enabling them to better process the relationship between the question and the images.
  • Inverted dynamics of open-sourced and proprietary models. Open-source models excel with a single clue image but fail to differentiate the clue when irrelevant images are also introduced; while proprietary models are initially conservative with a single clue but ultimately achieve better performance through multi-image augmentation. This reveals open-source models are weaker at handling multiple images, compared to proprietary models.
@@ 2 Related Work
While numerous benchmarks evaluate the capabilities of MLLMs, we focus here on those designed for knowledge-intensive visual question answering (VQA), where RAG is the primary solution paradigm. Table 1 compares the modalities covered by commonly used benchmarks. A more detailed discussion of how Visual-RAG differs from prior work is provided in Appendix A.
Textual Knowledge Centric VQA Benchmarks
OK-VQA (Marino et al., 2019) and its augmented version A-OKVQA (Schwenk et al., 2022) focus on commonsense or world knowledge, though they do not strongly depend on external knowledge retrieval. ViQuAE (Lerner et al., 2022), InfoSeek (Chen et al., 2023) and Encyclopedic-VQA (E-VQA, Mensink et al., 2023) address “encyclopedic” questions, paired with query images of specific entities, utilizing information retrieved from external textual sources (e.g., Wikipedia) to answer.
Visual Knowledge Centric QA Benchmarks
While the above benchmarks all position themselves as knowledge-intensive VQA tasks, it is important to note that the required knowledge remains primarily textual, with images functioning mostly as entity anchors. In contrast, to our knowledge, only two related works truly require image retrieval as evidence for answering questions.
WebQA (Chang et al., 2022) is an earlier effort that requires retrieving images for visual knowledge. However, it heavily relies on captions because the images themselves usually lack indicators of which are clue images. A concurrent work, MRAG-Bench (Hu et al., 2024) shares a similar motivation in using image retrieval for question answering. However, it primarily aims to recognize an entity when the instance in query image is incomplete or partially obscured, relying on retrieved images that present a complete view of the same entity. This entity-recognition objective differs from our focus on answering visual-feature-oriented factual queries. As queries in MRAG-Bench include image, the model can use the “shortcut” of matching a visually similar retrieved candidate, bypassing the need for deeper text-to-image reasoning.
@@ 3 Benchmark Overview
We propose Visual-RAG, a text-only Question Answering benchmark for challenging, visual knowledge intensive factual questions. It is an evaluation benchmark for visual knowledge centric text-to-image retrieval, as well as a benchmark for evaluating MLLMs’ abilities on extracting visual knowledge as augmentation under retrieval-augmented generation settings.
Visual Knowledge Intensive Queries
Our Visual-RAG focuses on text-only queries that probe factual knowledge of visual feature which generally hold true of an entity (e.g. “What colour are the stamens of Chamaecrista nictitans”), rather than instance-specific attributes commonly addressed in standard VQA tasks (e.g. “What colour is the cup on table”). Hence, we specifically select the organism domain which enables specialized queries. Unlike existing works, we deliberately exclude images from the queries due to the reason that including an image often reduces the problem to mere entity recognition or unimodal-similarity match, instead of our main objective of cross-modal knowledge extraction. Numerous established benchmarks already cover entity recognition, including the OVEN (Hu et al., 2023) dataset which serves as the foundation for the the InfoSeek dataset, and the iNaturalist 2021 (Van Horn et al., 2021) dataset, which we employ in this work. Moreover, text-only queries better align with real-world scenarios where users tend to raise questions about an entity without an image.
Naturally Co-occurring Hard Negative Images
For each question, multiple clue images might be available, reflecting real-world scenarios. Additionally, the dataset incorporates a substantial number of naturally co-occurring hard negative images, i.e., images of the same species that do not exhibit the visual feature in query. Notably, all clue images for a given question can be retrieved purely based on visual information, without reliance on captions.
Dataset Statistics
The current version of Visual-RAG contains 400 queries, forming an image knowledge base of total 103824 images. The average clue image count for each query is 13.89, and the clue images occupy 5.35% of the image knowledge base. The distribution of organism categories and query types is shown in Figure 2.
(a) Distribution of organism categories in Visual-RAG
(b) Distribution of question categories
Figure 2: Distribution of organisms and question categories.
@@ 4.1 Data source
Following Encyclopedic VQA Mensink et al. (2023), we adopt iNaturalist 2021 (iNat21, Van Horn et al., 2021) as the cornerstone of our benchmark. iNat21 contains 2.7M images of organism (animals, plants, fungi, etc.), spanning across 10000 taxonomy species. Each species in iNat21 has 200–300 images, originally collected for species-level image classification.
For organism species selection in query construction, we aim to ensure that the chosen species are relatively underrepresented in textual resources. To achieve that, we filter them based on the length of their Wikipedia summary and description sections. Well-known species typically have multiple sections (e.g., habitats, behaviour, etc.), indicating extensive documentation. We therefore include only those species whose summary and description account for no more than 50% length of the entire Wikipedia article. This also helps minimize reliance on prior knowledge: widely recognized species may have detailed descriptions on visual features across LLM pre-training corpora, even if these details are not explicitly mentioned in Wikipedia.

### v2

@@ Abstract
Retrieval‑augmented generation (RAG) is a paradigm that augments large language models (LLMs) with external knowledge to tackle knowledge‑intensive question answering. While several benchmarks evaluate Multimodal LLMs (MLLMs) under Multimodal RAG settings, they predominantly retrieve from textual corpora and do not explicitly assess how models exploit visual evidence during generation. Consequently, there still lacks benchmark that isolates and measures the contribution of retrieved images in RAG. We introduce Visual-RAG, a question‑answering benchmark that targets visually grounded, knowledge‑intensive questions. Unlike prior work, Visual-RAG requires text‑to‑image retrieval and the integration of retrieved clue images to extract visual knowledge necessary for answer generation. With Visual-RAG, we evaluate 5 open-source and 3 proprietary MLLMs and demonstrate that images provide strong evidence in augmented generation; however, even state‑of‑the‑art models struggle to effectively extract and utilize visual knowledge. Our results highlight the need for improved visual retrieval, grounding, and attribution in multimodal RAG systems. [^1] Benchmark available at: github.com/visual-rag/visual-rag
@@ 1 Introduction
Figure 1: InfoSeek style: Textual-Knowledge-Centric RAG (left), MRAG-Bench style: Retrieving common images for entity recognition and entity bridged prior knowledge recalling (middle) v.s. Ours: Visual Evidence-Centric Retrieval and Visual Knowledge-Intensive Reasoning (right). MRAG-Bench (Hu et al. 2025) formulates its task by presenting uncommon images that are challenging for entity recognition, retrieving more common images to support the MLLMs in bridging object identification. Typically, the actual answers (e.g., identifying the engine used in a specific car) is rarely explicitly represented visually in the retrieved images, making the answer dependent heavily on the model’s internal prior knowledge. In contrast, our Visual-RAG benchmark utilizes text-only queries with a common identifier (e.g., name of an organism), and requires retrieval of specialized, less common images containing explicit answer within their visual contents. Consequently, Visual-RAG makes cross-modal retrieval more challenging and demands deeper visually grounded knowledge acquisition, thus facilitating a more precise visual-centric evaluation of MLLMs’ capability to extract detailed visual information for knowledge-intensive reasoning.
Retrieval‑Augmented Generation (RAG) augments large language models (LLMs) with external knowledge, substantially improving performance on knowledge‑intensive queries (Lewis et al. 2020; Karpukhin et al. 2020; Izacard and Grave 2021). While the application of RAG is well-acknowledged for text-only knowledge-intensive tasks, many real‑world scenarios require multimodal knowledge from visual cues, such as diagrams or maps, highlighting the limitations of relying solely on text retrieval. Several benchmarks have been proposed to evaluate Multimodal Large Language Models (MLLMs) under multimodal RAG settings, targeting complex, knowledge‑intensive questions. Notably, InfoSeek (Chen et al. 2023) and Encyclopedic-VQA (Mensink et al. 2023) pair mixed‑modal queries directly with images. Yet, these benchmarks are fundamentally Textual‑Knowledge‑Centric: answering typically involves first recognizing entities in the query image and then retrieving textual evidence (e.g., Wikipedia passages) to generate the answer, as demonstrated in Figure 1 (left). As a result, although images are present in query, MLLMs rarely need to extract visual knowledge from them during answer generation, and current multimodal RAG benchmarks predominantly assess textual rather than visual knowledge exploitation.
Recent work MRAG-Bench (Hu et al. 2025) partially addresses this gap by explicitly requiring image retrieval. As shown in Figure 1 (middle), MRAG-Bench features challenging images as queries (e.g., damaged, partially occluded or incomplete entities), enforcing models to retrieve more common images of the entity to assist entity recognition. However, final answers (e.g., the engine used in a specific car model) largely rely on parametric model knowledge rather than explicit visual evidence contained in the retrieved images. Consequently, solving the task still depends on internal priors instead of information extracted from the pictures, leaving the core challenge of visually grounded RAG unaddressed.
We therefore introduce Visual-RAG, a visual knowledge‑intensive QA benchmark that explicitly requires text‑to‑image retrieval and visually grounded answer generation. Crucially, the retrieved image itself serves as evidence, which aligns with real-world use case: once a clue image containing the relevant visual feature is surfaced, one can answer the visual knowledge-intensive question directly, without treating the image as a pivot to trigger subsequent recalling of prior knowledge. Accordingly, Visual-RAG emphasizes Visual Evidence‑Centric Retrieval and mandates Visual Knowledge‑Intensive Reasoning, challenging models to extract and ground the information contained in clue images, and to discriminate it from visually similar hard-negative distractors. Key features of Visual-RAG include:
@@
    Fine‑grained and less common visual evidence. Only 6.12% of corpus images are clue images containing the necessary visual evidence, substantially making cross‑modal retrieval challenging.
@@
    Hard negatives at scale. Most non‑clue images are hard negatives, which depict the same entities and are visually similar, stress‑testing both retrieval models’ robustness and downstream MLLMs’ capabilities in visual knowledge extraction.
@@
    Explicit visual answerability. Evidences of answers are encoded directly in the visual content, enabling evaluation of whether MLLMs truly leverage visual information.
To construct Visual-RAG, we initially generate candidate queries using LLMs, followed by rigorous refinement through human filtering and rewriting processes. Subsequently, we employ both open-source MLLMs and human annotators for additional filtering, ensuring the final queries reflect genuine human curiosity and practical relevance.
Benchmark Query Modality Knowledge Modality Retrieval Modality
InfoSeek (Chen et al. 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
E-VQA (Mensink et al. 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
WebQA (Chang et al. 2022) [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
MRAG-Bench (Hu et al. 2025) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Visual-RAG (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Table 1: Comparing modalities of knowledge-intensive VQA benchmarks. The link icon [Uncaptioned image] denotes that the two objects are paired, retrieving one will automatically link to the other. For InfoSeek and E-VQA, Yan and Xie (2024) demonstrated that by image-to-image retrieval on Wikipedia images, the Wikipedia articles containing those images are also retrieved at a significantly higher recall comparing to image-to-text retrieval. Similarly for WebQA, a question-to-caption (text-to-text) retrieval achieves better performance than text-to-image retrieval, as shown in Appendix A. Our Visual-RAG enforces text-to-image retrieval for visual knowledge.
With Visual-RAG, we comprehensively evaluate eight popular MLLMs, including five open-source and three proprietary models. Our baseline results confirm that images can serve as strong evidence for augmented generation. We also examine multiple RAG configurations to assess how effectively MLLMs ingest image augmentation, and to analyze their ability to discriminate informative visual clues from irrelevant distractors. The key findings from our experiments are summarized as follows:
@@
  • Visual evidence-centric retrieval is challenging. The commonly used cross-modal retrievers face significant challenge in addressing difficult queries requiring the identification of fine-grained visual evidence.
@@
  • Visual clues can serve as effective evidence for Retrieval-Augmented Generation. The evaluated MLLMs demonstrate the ability to extract visual knowledge to answer questions when presented with single ground-truth clue image.
  • Realistic visual RAG remains an onerous task. With top-$k$ RAG, performance of most models lags behind the ground-truth clue upper bound. Proprietary models reach the upper bound only when $k$ close to 20, revealing limited ability to efficiently exploit visual clues, partially constrained by the bottleneck of retriever.
  • Inverted dynamics of open-source and proprietary models. When a guaranteed clue is mixed with non-clue images, open-source MLLMs leverage it well at lower $k$, but their accuracy decays as including more distractors. In contrast, proprietary MLLMs retain robustness in identifying visual evidence from distractors at higher $k$, yet overlook the clue when only a few images are provided.
@@ 2 Related Work
Numerous benchmarks have been developed to evaluate the capabilities of MLLMs. We focus on those specifically designed for knowledge-intensive visual question answering (VQA), particularly emphasizing retrieval-augmented generation (RAG) paradigms. Table 1 provides an overview and comparison of modalities across widely adopted benchmarks.
Textual Knowledge Centric VQA Benchmarks.
Benchmarks such as OK-VQA (Marino et al. 2019) and its augmented variant, A-OKVQA (Schwenk et al. 2022), predominantly emphasize commonsense and world knowledge but exhibit limited dependence on external knowledge retrieval mechanisms. Other datasets like ViQuAE (Lerner et al. 2022), InfoSeek (Chen et al. 2023) and Encyclopedic-VQA (E-VQA, Mensink et al. 2023) address encyclopedic questions paired with images of specific entities. However, the retrieved information employed in answering is primarily textual (e.g., passages from Wikipedia), with the visual modality mainly serving as an entity anchor rather than as a source of knowledge.
Visual Knowledge Centric QA Benchmarks.
To date, only a limited number of benchmarks explicitly require visual knowledge retrieval as evidence in answering questions. WebQA (Chang et al. 2022) is an earlier effort that necessitates retrieving images for visual knowledge but relies heavily on textual captions, as images in the benchmark often lack explicit indicators for being clue. More recently, MRAG-Bench (Hu et al. 2025) similarly emphasizes image retrieval. However, MRAG-Bench provides queries in the form of challenging images depicting partially obscured or incomplete entities, aiming primarily at entity recognition through matching visually similar images. The retrieved images predominantly help in identifying entities, with answers typically dependent on the model’s prior knowledge rather than directly on visual evidence within these images.
Unlike MRAG-Bench, our benchmark, Visual-RAG , explicitly constructs textual queries and requires retrieving specialized images that explicitly contain visual evidence. Consequently, our benchmark evaluates deeper visual knowledge extraction capabilities for MLLMs, emphasizing visual evidence-centric reasoning. More detailed comparison between Visual-RAG and prior benchmarks is provided in Appendix A.
@@ 3 Benchmark Overview
We propose Visual-RAG, a Question Answering (QA) benchmark for visual knowledge–intensive factual queries. Visual-RAG evaluates two capabilities under retrieval‑augmented generation (RAG): (i) visual evidence–centric text‑to‑image retrieval and (ii) an MLLM’s ability for visual knowledge-intensive reasoning with retrieved images in retrieval-augmented answer generation.
Task Definition.
Let $\mathcal{Q}$ be a set of text queries and $\mathcal{D}$ a corpus of images. For each $q\in\mathcal{Q}$ there exists a small set of clue images $\mathcal{I}_{c}\subset\mathcal{D}$ whose pixel content explicitly encodes the visual evidence needed to answer $q$. The remaining images $\mathcal{I}_{d}=\mathcal{D}\setminus\mathcal{I}_{c}$ act as distractors; many are hard negatives that depict the same entity or highly similar instances but lack the queried visual feature.
Given $q$, a system first retrieves $k$ images $\mathcal{I}_{ret;k}\subset\mathcal{D}$ and then generates an answer $\hat{y}$ conditioned on $(q,I_{ret;k})$. Crucially, Visual-RAG assumes that once a correct clue image is retrieved, the answer is directly derivable from that image, without requiring subsequent textual lookup.
Visual Knowledge–Intensive Queries.
Queries in Visual-RAG are text‑only. They probe factual properties of visual features that generally hold for an entity category (e.g., “What color are the hindwings of Papaipema inquaesita?”, as in Figure 1), rather than instance‑specific properties typical of standard VQA (e.g., “What color is the cup on the table?”). We focus on the organism domain because it naturally supports fine‑grained, specialized questions for which the necessary evidence is visual and stable at the category level. Unlike many existing works, we deliberately exclude image from the query input to avoid reducing the task to entity recognition or image-to-image unimodal similarity matching; the primary objective is cross‑modal knowledge extraction from retrieved images. Numerous established benchmarks already cover entity recognition, including the OVEN (Hu et al. 2023) dataset which serves as the foundation for the InfoSeek dataset, and the iNaturalist 2021 (Van Horn et al. 2021) dataset, which we employ in this work. Moreover, our setting of visual knowledge-intensive queries simulates realistic application scenarios, where referring directly to images for answers is more intuitive and efficient compared to searching through textual documentation for such query.
Naturally Co-occurring Hard Negative Images.
Visual-RAG incorporates a large number of naturally co‑occurring hard negatives: images of the same species that do not exhibit the queried feature. All clue images are retrievable based on visual content alone; captions or alt‑text are not required or assumed at retrieval time. This construction jointly pressures the retriever - to surface the rare, feature‑bearing images, and the answer generator MLLM - to extract and ground the relevant visual evidence.
Dataset Statistics.
Our Visual-RAG contains 374 queries, forming an image knowledge base of total 99,017 images. On average, each query has 16.19 clue images; across the corpus, clue images constitute 6.12% of all images. The distributions of organism categories and query types are shown in Figure 2.
(a) Organism categories.
(b) Question categories.
Figure 2: Distribution of organisms and question categories in Visual-RAG .
@@ 4.1 Data source
Following Encyclopedic VQA (Mensink et al. 2023), we adopt iNaturalist 2021 (iNat21, Van Horn et al. 2021) as the cornerstone of our benchmark. iNat21 contains 2.7M images of organism (animals, plants, fungi, etc.), spanning across 10000 taxonomy species. Each species in iNat21 has 200–300 images, originally collected for species-level image classification.